PDS_VERSION = PDS3 RECORD_TYPE = STREAM OBJECT = TEXT PUBLICATION_DATE = 1997-02-13 NOTE = "MGS RST Instrument Health Report HEA7044A.TXT" END_OBJECT = TEXT END Engineering Data ================ Engineering Data Statistical Summary from CSV2STAT: First SCET: 1997-02-12T00:50:08.755 Last SCET: 1997-02-12T23:59:59.751 CHANNL CHANNEL TITLE DN-LO DN-HI EU-LO EU-HI EU-AVG/STDEV RECDS ------ ------------- ----- ----- --------- --------- ---------/-------- ----- L-0037 MOT1_USO_ENA 0 0 35362 L-0047 MOT2_USO_ENA 1 1 35362 L-0060 TWTA1_FILMNT 0 0 35362 L-0061 TWTA1_HV 0 0 35375 L-0070 TWTA2_FILMNT 1 1 35375 L-0071 TWTA2_HV 0 0 35362 L-0080 USO_PWR 0 0 35362 L-0131 TWTA1_HLX_I 0 1 -3.84E-02 -1.85E-02 -3.83E-02/8.77E-06 17792 L-0132 TWTA1_ANOD_V 0 1 -1.00E+03 -1.00E+03 -1.00E+03/1.90E-03 17599 L-0141 TWTA2_HLX_I 38 43 7.05E-01 8.05E-01 7.56E-01/1.12E-04 17792 L-0142 TWTA2_ANOD_V 59 64 -8.70E+02 -8.60E+02 -8.65E+02/1.19E-02 17599 L-0200 USO_REG_V 203 205 4.06E+00 4.10E+00 4.08E+00/5.51E-05 17599 L-0201 USO_OVEN_V 123 126 2.46E+00 2.52E+00 2.48E+00/7.43E-05 17599 T-0302 EPC1_T 105 126 2.50E+01 3.41E+01 2.85E+01/3.44E-02 4420 T-0303 EPC2_T 100 130 2.33E+01 3.63E+01 2.70E+01/5.44E-02 4420 T-0310 RF_ISOLTR1_T 171 200 2.03E+01 3.14E+01 2.36E+01/2.23E-02 17793 T-0311 RF_ISOLTR2_T 157 180 2.82E+01 3.63E+01 3.08E+01/1.61E-02 17793 T-0312 TWT1_T 170 201 1.98E+01 3.17E+01 2.31E+01/4.89E-02 4420 T-0313 TWT2_T 152 177 2.93E+01 3.82E+01 3.19E+01/3.60E-02 4420 T-0315 USO_T 176 180 9.93E+00 1.12E+01 1.07E+01/2.39E-03 4420 L-0200 remains constant at 4.08 volts. L-0201 has dropped back from 2.49 to 2.48 volts (average). T-0315 has dropped back from 11.1C to 10.7C (it was 10.6C two days ago). Open Loop Data ============== Data Collection --------------- Trish Priest reported that USO Test #6 looked good, after some problems with attenuator settings were ironed out. We queried for data over less than the full interval; our results were: YYYY/DDD DSS DSP SAMP BW DSP BIT TRK REC RECS RECS FIRST START STOP RATE MODE RESN MODE LEN MSSG SNR -------- --- -------- -------- ---- ---- ---- ---- ---- ---- ----- ---- ----- 1997/043 45 13:00:00 14:46:00 5000 2000 2 12 1-W 1666 31794 7 ~57. Data Anomalies -------------- The query was issued after a retransmission had been completed. In addition to the 7 missing records, there was one duplicate record. This is the first time we have had missing records following a retransmission. Frequency Analyses ------------------ Joe Twicken has completed his analysis of the data. His report: I have processed the ODS data from 97/043. This ODS was recorded at DSS 45. The data were acquired as part of USO Test #6. The SNR was approximately 59 dB/Hz during the Test. This was the third USO test since the MGS HGA was deployed. The ODS was acquired at DSS 45 from 13:00:00 to 14:46:00 at 5000 samples per second in mode 2 (1250 12-bit samples/second for each of 4 A/D's). The USO test itself began at 12:45:00, but there were a number of attenuator changes at the beginning of the test so we did not query for the first 15 minutes worth of data. There were also a couple of attenuator changes during the period of data that I processed. I used the spk_c_961111-970328 SP-kernel file to steer the data. This file was created on 97/029. The ODS data was retransmitted once from the DSN station. Nevertheless, there were 7 missing ODS records and 1 duplicate ODS record. If all records had been available, we would have received 31800 records for the period of our query. The missing records after a retransmission of the ODS are disappointing. The POCA file was fixed with our "salvagePoca" program. No hand modification of the POCA file was required. A very brief analysis of the original POCA file indicated that approximately 100 POCA records were in error. During the 1 hour and 46 minute recording there were 6360 POCA records (1 per second). An amplitude modulation with a period of about 5+ minutes is evident in all the carrier amplitude estimates. The peak-to-peak amplitude variation during this test was far less than that observed in USO Test #4 and the period was somewhat shorter. In this test, the peak-to-peak variation was on the order of 0.2 to 0.3 dB. The source of this modulation remains unknown. As in some previous tests, there were also a pair of frequency spurs at approximately 650 Hz on either side of the carrier. In this test, the level of these spurs was about 32 to 35 dB below that of the carrier. The spurs appear to drift through the spectrum along with the carrier. I estimated the USO frequency to be 8423152923.15 Hz. The best linear fit to the USO frequency since 96/323 has a slope of 0.249 Hz/day. This fit is actually quite good. The 1-sigma error between our actual USO frequency estimates since 96/323 and the best linear fit to them is 0.32 Hz. The Allan deviation for this recording is shown below. The SNR during the acquisition of data was approximately 59 dB/Hz. I did not compensate for the spacecraft spin cycle. The Allan deviation at the shorter integration times is overestimated due to the 7 missing data records. It was not even possible to make a decent estimate at an integration time of 0.4 seconds. The estimates for 97/043 are: t (sec) AD ------ ---------- 1.6384 4.0302e-13 3.2768 2.8365e-13 6.5536 2.3127e-13 26.2144 1.6758e-13 I would like to point out that I have put our MGS database ("MgsDB") into operation this week, and have begun recording RS and auxiliary data files with the data base. All files transfered to Stanford for RS purposes either from the CDB, the NERT cache or by FTP will be recorded in the MgsDB together with the PDS label keywords and values generated for those files. In addition, all processed open-loop data files will be recorded in the MgsDB along with the program name and version used to create them, the time that the program was executed, and the values of all relevant input parameters/file names used for creation of the files. The MgsDB has both a GUI and a scriptable interface. Closed Loop Data ================ We obtained seven ODF's from OSCAR yesterday. None have been examined. 97034C035.DAT;1 97035C036.DAT;1 97036C037.DAT;1 97037C038.DAT;1 97038C041.DAT;1 97041C042.DAT;2 97042C043.DAT;1 We looked for but did not find UTPM and WEATHER data on OSCAR. We obtained the two most recent *.short files from the EOP Web site, but were unable to complete transfer of the corresponding two *.long files. We have submitted material for an ISA to Trish.